# User Experience in Dataset Search Platform Interfaces

### v1

@@ Abstract.
This research investigates User Experience (UX) issues in dataset search platform interfaces, targeting Google Dataset Search and data.europa.eu. It focuses on 6 areas within UX: Initial Interaction, Search Process, Dataset Exploration, Filtering and Sorting, Dataset Actions, and Assistance and Feedback. The evaluation method combines ’The Pandemic Puzzle’ user task, think-aloud methods, and demographic and post-task questionnaires. 29 strengths and 63 weaknesses were collected from 19 participants involved in roles within technology firm or academia. While certain insights are specific to particular platforms, most are derived from features commonly observed in dataset search platforms across a variety of fields, implying that our findings are broadly applicable. Observations from commonly found features in dataset search platforms across various fields have led to the development of 10 new design prototypes. Unlike literature retrieval, dataset retrieval involves a significant focus on metadata accessibility and quality, each element of which can impact decision-making. To address issues like reading fatigue from metadata presentation, inefficient methods for results searching, filtering, and selection, along with other unresolved user-centric issues on current platforms. These prototypes concentrate on enhancing metadata-related features. They include a redesigned homepage, an improved search bar, better sorting options, an enhanced search result display, a metadata comparison tool, and a navigation guide. Our aim is to improve usability for a wide range of users, including both developers and researchers.
@@
User study, User interaction, Interview, Think-aloud, Search interface, Search platform, Dataset search, Dataset retrieval, Dataset, Metadata, Structured data, Information search, Information retrieval, Google dataset search, data.europa.eu
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; 2024; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Information retrievalccs: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing User studiesccs: Information systems Search interfacesccs: Human-centered computing Accessibilityccs: Human-centered computing Empirical studies in HCIccs: Information systems Web searching and information discovery
1. Introduction
@@
In the present data-driven landscape, sectors such as healthcare and social sciences are increasingly reliant on diverse datasets (Benjelloun et al., 2020; Mayer-Schönberger and Cukier, 2013). The sheer volume and complexity of available datasets often result in data silos, hampering their effective usage (Roh et al., 2019; Stonebraker et al., 2018). This dichotomy between the surfeit of publicly available datasets and the particular datasets users can actually locate constitutes a significant challenge (Chapman et al., 2020).
@@
Earlier studies have delved into the intricacies of dataset search systems, highlighting the critical role of metadata and the need for its standardisation (Chapman et al., 2020; Benjelloun et al., 2020). Research has also been conducted on how Google Dataset Search leverages metadata to organise and rank search results (Brickley et al., 2019). Additionally, some focus has been given to understanding user behaviour on governmental dataset platforms, identifying the primary factors that influence search queries (Kacprzak et al., 2019). However, there still exists a void in understanding the user experience in navigating these dataset search platforms, an area this research aims to explore.
@@
For this investigation, Google Dataset Search and data.europa.eu were selected due to their advanced metadata-based search features and high prevalence among researchers. This study assesses user experience in both document retrieval and dataset retrieval, examining existing features of Google Dataset Search and data.europa.eu platforms. Researchers synthesis these elements to pinpoint six aspects of user experience in dataset search. A user task, termed ”The Pandemic Puzzle,” is then designed to explore currently trending areas in dataset search. This study adopts both Concurrent and Stimulated Retrospective think-aloud techniques to investigate user interactions with these dataset retrieval platforms. To enhance the reliability of the think-aloud methods, Thought Prompts and preliminary training sessions are included. Additional context and a swift assessment of perceived usability are provided through demographic details and post-task questionnaires.
@@ 2. Related Works
Information seeking behaviour: in web searches and digital libraries, information seeking behaviour is a well-studied area with established models that map out the process and strategies users follow  (Ellis, 1989). However, the suitability of these models for dataset retrieval is uncertain since they were originally intended for document searches  (Ellis, 1989). Searches are generally of two types: ’lookup’, which is simple and well-supported, and ’exploratory’, which is more complex, involving tasks that require users to learn and discover new information  (Marchionini, 2006). Exploratory searches can be open-ended, leading to uncertainty when information is incomplete  (Marchionini, 2006). The complexity of exploratory searches is especially significant in dataset retrieval, which has become more challenging with initiatives like GO FAIR that encourage the sharing and reuse of datasets. Such initiatives have led to a greater volume and variety of data, highlighting the need for dataset search systems to be designed with the user’s needs in mind, helping them navigate the complex process of finding the right datasets  (Kern and Mathiak, 2015).
@@
Nature of datasets: Chapman et al.  (Chapman et al., 2020) describe a dataset as an ordered collection of either numerical or textual data points. Metadata, crucial for search enhancement and contextualisation, includes elements like data source, format, and collection time. Their work identifies two core dataset search approaches: ’basic’ and ’user-organised.’ The ’basic’ search is found on platforms such as Figshare (Thelwall and Kousha, 2016) and Google Dataset Search (Brickley et al., 2019), where datasets are pre-curated and stored in repositories. The ’user-organised’ search takes place in environments like data lakes (Gao et al., 2018) and data markets (Grubenmann et al., 2018), where users assemble individual data points themselves, as seen in climate studies collecting temperature and humidity metrics in a data lake.
@@
Online Dataset Accessibility: Benjelloun et al. (Benjelloun et al., 2020) examined Schema.org  (sch, 2024), noted for having the largest and most varied corpus of its kind, and found that only 34% of datasets include licensing information, while 44% offer data download links. Additionally, metadata coverage generally falls under 50%, affecting dataset reusability and findability, thereby influencing user experience in dataset searches. Their study also highlights that 37% of datasets lack persistent URLs, underscoring the necessity for stable identifiers like DOIs. Their findings advocate for standardisation in dataset descriptions, licensing, and access. While transitional periods will be required for standardisation, immediate improvements in user experience are achievable.
@@
Mechanism for searching datasets: Brickley et al. (Brickley et al., 2019) describe the dataset search engine’s multi-stage process, starting with web crawlers collecting metadata across the web. This metadata is subsequently standardised and structured into a universal format using frameworks like Schema.org. Cleaned metadata is then matched with a pre-existing database, or a knowledge graph, to add contextual richness to each dataset. The engine assigns unique ’fingerprints’ to datasets to group similar datasets and eliminate redundancies. Brickley et al. also discuss the engine’s use of MapReduce (Brickley et al., 2019) for efficient data processing. Daily scans are performed to capture new or updated datasets. During a search, the engine ranks datasets based on their query relevance, continually updating its methods to improve its effectiveness and reach.
@@
User-centric literature retrieval: from literature  (Ghorab et al., 2013; Kelly et al., 2015; Croft et al., 2010), they summary 10 user considerations: ’Relevance of Results’ focuses on query interpretation accuracy, context sensitivity, and result ordering; ’Ease of Use’ highlights intuitive interfaces and straightforward navigation; ’Performance and Speed’ aims for quick responses and system reliability; ’Search Features’ includes advanced search options and auto-suggestions; ’Quality of Information’ stresses source credibility and data precision; ’Security and Privacy’ emphasizes user consent and encryption; ’Adaptability’ caters to different languages, regions, and devices; ’Presentation of Results’ targets clear labelling and data visualisation options; and ’Personalisation’ utilises user history and preferences for customisation. In addition, ’Accessibility features’, such as voice search and readability, are integrated to ensure inclusive design.
@@
User-centric dataset retrieval: from literature  (Chapman et al., 2020; Kacprzak et al., 2019; Benjelloun et al., 2020), they indicates that factors like search result accuracy, user interface clarity, search speed, data protection, and system adaptability are crucial . Four key areas for tool enhancement emerge: ’Dataset Metadata’ should have comprehensive descriptions, standardisation, frequent updates, and be optimised for search; ’Functionality, Support, & Security’ implies powerful search features, reliable support, robust data security, and wide accessibility; ’Usability’ entails a user-friendly interface, fast performance, and effective visualisation; ’Support’ is augmented by detailed documentation, educational resources, active user communities, and ongoing tool improvements.
@@
Comparing dataset retrieval with traditional literature retrieval: a study  (Kern and Mathiak, 2015) with 53 participants highlighted the necessity for dataset retrieval systems to be tailored to the specific needs of researchers, as their decision-making process is more complex than that for literature retrieval. The study’s key findings emphasized the significance of metadata quality and accessibility, with users showing a willingness to contribute to metadata enhancement. Additionally, the research found that users value supplementary literature for added context and prefer search systems that are intuitive, like those with autocomplete features, and that offer research-relevant updates and suggestions.
@@
User study on dataset search platform: Kacprzak et al. (Kacprzak et al., 2019) scrutinise user interactions on government dataset search platforms, employing both qualitative and quantitative methods. They pinpoint two significant factors influencing user queries: data features and structural layout. Their research indicates user queries often include geospatial and temporal elements. Metrics reveal that search activity peaks during typical working hours, likely indicating professional use. Their study’s data also show most users access platforms via search engines and predominantly use browsers like Chrome, although Internet Explorer retains significant usage, possibly due to its corporate prevalence. They observe that user engagement metrics, such as page views and time spent, increase with return visits, and they identify challenges users face in data discovery as a prospective area for enhancing search functionalities.
3. Approach
@@
3.1. Explore Platform Features
@@
3.1.1. Google Dataset Search
@@
Figure 1. Google Dataset Search Platform Features
Highlight each specific element of the Google Dataset Search interface with a red box.
Figure 1. Google Dataset Search Platform Features
3.1.2. data.europa.eu
@@
Figure 2. data.europa.eu Platform Features
Highlight each specific element of the data.europa.eu interface with a red box.
Figure 2. data.europa.eu Platform Features
3.2. Explore User Experience Aspects
@@
3.3. Design User Task
@@
Table 1. The Pandemic Puzzle User Task
@@
Background
As a health sciences researcher, investigate COVID-19’s impacts on the UK’s life and social sciences sectors. Focus on datasets that are current, freely accessible, commercial-use, well-documented, from reputable sources, and specific to UK regions in tabular form.
1. Initial Interaction
Engage with the platform to locate datasets on ’COVID-19 impact on UK healthcare systems’ and ’social science research during pandemics’. Assess interface usability including visibility of search, delete, and return functions.
2. Search Process
Input specific queries like ’COVID-19 mental health data UK’ and observe platform’s auto-complete suggestions, result speed, and relevance to the UK’s life and social sciences.
3. Dataset Exploration
Scrutinize datasets for relevance, examining names, sources, types, update frequency, geographic specificity, and alignment with the research on UK’s response to COVID-19.
4. Filters and Sorting
Utilize filters to narrow results to recent, open-access datasets, applying sorting to prioritize by relevance or recency. Assess the precision of filtering and sorting mechanisms.
5. Dataset Actions
Identify suitable datasets for longitudinal analysis of the pandemic’s effects, noting options for saving and sharing. Evaluate the efficiency and intuitiveness of these processes.
6. Feedback and Help
Document any usability issues encountered and identify how the platform provides assistance. Explore the process for submitting feedback and seeking help within the platform.
3.4. Use of User Study Methods
3.4.1. Use of Concurrent Think-Aloud
This study uses moderate-complexity tasks, guided by (Kelly et al., 2009). Participants perform tasks that mix quick reactions with thoughtful analysis, such as choosing a relevant dataset and analysing its metadata. These tasks are more involved than basic queries but not as complex as detailed data analysis. This research employs a think-aloud technique as suggested by (Charters, 2003), giving real-time insight into participants’ thinking. VooV software captures both audio and visual data for comprehensive documentation.
3.4.2. Challenges in Concurrent Think-Aloud
One issue is that participants may feel awkward vocalising their thoughts, which can affect natural decision-making and skew results. Participants might also forget to verbalise their thoughts, especially when focused or performing difficult tasks. Another challenge is the risk of cognitive overload due to multitasking, especially in tasks requiring intense concentration (Charters, 2003).
3.4.3. Addressing Challenges During Concurrent Think-Aloud
@@
Used to help participants articulate their thoughts during this study. These are specific questions or statements aimed at prompting real-time reactions as participants use dataset search platforms. The goal is to get a detailed understanding of the user experience. Examples of thought prompts include: ”What do you think of this dataset after selecting it?” and ”Which filters are you using and why?”.
3.4.4. Use of Stimulated Retrospective Think-Aloud
This research incorporates both Stimulated Retrospective Think-Aloud (SRTA) and Concurrent Think-Aloud (CTA) techniques to achieve a comprehensive view of user interactions (SELA et al., 2000; Charters, 2003). The SRTA method allows participants to reflect on their actions and thoughts, using visual aids or playback footage for reference. Structured SRTA questions guide this reflective process, creating a reliable setting for participants to share their experiences, thereby improving data quality. An example from this study illustrates this approach.
  • Immediate Thought Prompt: ”What are your first impressions of the search results? Are they what you expected?”
  • Possible Immediate Answer: ”There are more results than I anticipated, making it overwhelming. I expected fewer or more targeted outcomes.”
  • Retrospective Thought Prompt: ”Reflecting on the search, do you remember your initial reactions? Did the results meet your expectations?”
  • Possible Retrospective Answer: ”Initially, I was shocked by the sheer volume of information. I now realise my search terms were too broad, leading to the overload. If I were to do it again, I’d use more specific search criteria to narrow down the results.”

### v2

@@ Abstract
This research investigates User Experience (UX) issues in dataset search, targeting Google Dataset Search and data.europa.eu. It focuses on 6 areas within UX: Initial Interaction, Search Process, Dataset Exploration, Filtering and Sorting, Dataset Actions, and Assistance and Feedback. The evaluation method combines ’The Pandemic Puzzle’ user task, think-aloud methods, and demographic and post-task questionnaires. 29 strengths and 63 weaknesses were collected from 19 participants involved in roles within technology firm or academia. While certain insights are specific to particular platforms, most are derived from features commonly observed in dataset search platforms across a variety of fields, implying that our findings are broadly applicable. Observations from commonly found features in dataset search platforms across various fields have led to the development of 10 new design prototypes. Unlike literature retrieval, dataset retrieval involves a significant focus on metadata accessibility and quality, each element of which can impact decision-making. To address issues like reading fatigue from metadata presentation, inefficient methods for results searching, filtering, and selection, along with other unresolved user-centric issues on current platforms. These prototypes concentrate on enhancing metadata-related features. They include a redesigned homepage, an improved search bar, better sorting options, an enhanced search result display, a metadata comparison tool, and a navigation guide. Our aim is to improve usability for a wide range of users, including both developers and researchers.
Keywords:
@@
User study User interaction Interview Think-aloud Search interface Search platform Dataset search Dataset retrieval Dataset Metadata Structured data Information search Information retrieval Google dataset search data.europa.eu
1 Introduction
@@
In the present data-driven landscape, sectors such as healthcare and social sciences are increasingly reliant on diverse datasets [3, 20]. The sheer volume and complexity of available datasets often result in data silos, hampering their effective usage [21, 23]. This dichotomy between the surfeit of publicly available datasets and the particular datasets users can actually locate constitutes a significant challenge [6].
@@
Earlier studies have delved into the intricacies of dataset search systems, highlighting the critical role of metadata and the need for its standardisation [6, 3]. Research has also been conducted on how Google Dataset Search leverages metadata to organise and rank search results [5]. Additionally, some focus has been given to understanding user behaviour on governmental dataset platforms, identifying the primary factors that influence search queries [14]. However, there still exists a void in understanding the user experience in navigating these dataset search platforms, an area this research aims to explore.
@@
For this investigation, Google Dataset Search and data.europa.eu were selected due to their advanced metadata-based search features and high prevalence among researchers. This study assesses user experience in both document retrieval and dataset retrieval, examining existing features of Google Dataset Search and data.europa.eu platforms. Researchers synthesis these elements to pinpoint six aspects of user experience in dataset search. A user task, termed "The Pandemic Puzzle," is then designed to explore currently trending areas in dataset search. This study adopts both Concurrent and Stimulated Retrospective think-aloud techniques to investigate user interactions with these dataset retrieval platforms. To enhance the reliability of the think-aloud methods, Thought Prompts and preliminary training sessions are included. Additional context and a swift assessment of perceived usability are provided through demographic details and post-task questionnaires.
@@ 2 Related Works
Information seeking behaviour: in web searches and digital libraries, information seeking behaviour is a well-studied area with established models that map out the process and strategies users follow  [9]. However, the suitability of these models for dataset retrieval is uncertain since they were originally intended for document searches  [9]. Searches are generally of two types: ’lookup’, which is simple and well-supported, and ’exploratory’, which is more complex, involving tasks that require users to learn and discover new information  [19]. Exploratory searches can be open-ended, leading to uncertainty when information is incomplete  [19]. The complexity of exploratory searches is especially significant in dataset retrieval, which has become more challenging with initiatives like GO FAIR that encourage the sharing and reuse of datasets. Such initiatives have led to a greater volume and variety of data, highlighting the need for dataset search systems to be designed with the user’s needs in mind, helping them navigate the complex process of finding the right datasets  [17].
@@
Nature of datasets: Chapman et al.  [6] describe a dataset as an ordered collection of either numerical or textual data points. Metadata, crucial for search enhancement and contextualisation, includes elements like data source, format, and collection time. Their work identifies two core dataset search approaches: ’basic’ and ’user-organised.’ The ’basic’ search is found on platforms such as Figshare [24] and Google Dataset Search [5], where datasets are pre-curated and stored in repositories. The ’user-organised’ search takes place in environments like data lakes [10] and data markets [12], where users assemble individual data points themselves, as seen in climate studies collecting temperature and humidity metrics in a data lake.
@@
Online Dataset Accessibility: Benjelloun et al. [3] examined Schema.org  [2], noted for having the largest and most varied corpus of its kind, and found that only 34% of datasets include licensing information, while 44% offer data download links. Additionally, metadata coverage generally falls under 50%, affecting dataset reusability and findability, thereby influencing user experience in dataset searches. Their study also highlights that 37% of datasets lack persistent URLs, underscoring the necessity for stable identifiers like DOIs. Their findings advocate for standardisation in dataset descriptions, licensing, and access. While transitional periods will be required for standardisation, immediate improvements in user experience are achievable.
@@
Mechanism for searching datasets: Brickley et al. [5] describe the dataset search engine’s multi-stage process, starting with web crawlers collecting metadata across the web. This metadata is subsequently standardised and structured into a universal format using frameworks like Schema.org. Cleaned metadata is then matched with a pre-existing database, or a knowledge graph, to add contextual richness to each dataset. The engine assigns unique ’fingerprints’ to datasets to group similar datasets and eliminate redundancies. Brickley et al. also discuss the engine’s use of MapReduce [5] for efficient data processing. Daily scans are performed to capture new or updated datasets. During a search, the engine ranks datasets based on their query relevance, continually updating its methods to improve its effectiveness and reach.
@@
User-centric literature retrieval: from literature  [11, 15, 8], they summary 10 user considerations: ’Relevance of Results’ focuses on query interpretation accuracy, context sensitivity, and result ordering; ’Ease of Use’ highlights intuitive interfaces and straightforward navigation; ’Performance and Speed’ aims for quick responses and system reliability; ’Search Features’ includes advanced search options and auto-suggestions; ’Quality of Information’ stresses source credibility and data precision; ’Security and Privacy’ emphasizes user consent and encryption; ’Adaptability’ caters to different languages, regions, and devices; ’Presentation of Results’ targets clear labelling and data visualisation options; and ’Personalisation’ utilises user history and preferences for customisation. In addition, ’Accessibility features’, such as voice search and readability, are integrated to ensure inclusive design.
@@
User-centric dataset retrieval: from literature  [6, 14, 3], they indicates that factors like search result accuracy, user interface clarity, search speed, data protection, and system adaptability are crucial . Four key areas for tool enhancement emerge: ’Dataset Metadata’ should have comprehensive descriptions, standardisation, frequent updates, and be optimised for search; ’Functionality, Support, & Security’ implies powerful search features, reliable support, robust data security, and wide accessibility; ’Usability’ entails a user-friendly interface, fast performance, and effective visualisation; ’Support’ is augmented by detailed documentation, educational resources, active user communities, and ongoing tool improvements.
@@
Comparing dataset retrieval with traditional literature retrieval: a study  [17] with 53 participants highlighted the necessity for dataset retrieval systems to be tailored to the specific needs of researchers, as their decision-making process is more complex than that for literature retrieval. The study’s key findings emphasized the significance of metadata quality and accessibility, with users showing a willingness to contribute to metadata enhancement. Additionally, the research found that users value supplementary literature for added context and prefer search systems that are intuitive, like those with autocomplete features, and that offer research-relevant updates and suggestions.
@@
User study on dataset search platform: Kacprzak et al. [14] scrutinise user interactions on government dataset search platforms, employing both qualitative and quantitative methods. They pinpoint two significant factors influencing user queries: data features and structural layout. Their research indicates user queries often include geospatial and temporal elements. Metrics reveal that search activity peaks during typical working hours, likely indicating professional use. Their study’s data also show most users access platforms via search engines and predominantly use browsers like Chrome, although Internet Explorer retains significant usage, possibly due to its corporate prevalence. They observe that user engagement metrics, such as page views and time spent, increase with return visits, and they identify challenges users face in data discovery as a prospective area for enhancing search functionalities.
3 Approach
@@
3.1 Explore Platform Features
@@
3.1.1 Google Dataset Search
@@
Figure 1: Google Dataset Search Platform Features
3.1.2 data.europa.eu
@@
Figure 2: data.europa.eu Platform Features
3.2 Explore User Experience Aspects
@@
3.3 Design User Task
@@
Table 1: The Pandemic Puzzle User Task
@@
Background As a health sciences researcher, investigate COVID-19’s impacts on the UK’s life and social sciences sectors. Focus on datasets that are current, freely accessible, commercial-use, well-documented, from reputable sources, and specific to UK regions in tabular form.
1. Initial Interaction Engage with the platform to locate datasets on ’COVID-19 impact on UK healthcare systems’ and ’social science research during pandemics’. Assess interface usability including visibility of search, delete, and return functions.
2. Search Process Input specific queries like ’COVID-19 mental health data UK’ and observe platform’s auto-complete suggestions, result speed, and relevance to the UK’s life and social sciences.
3. Dataset Exploration Scrutinize datasets for relevance, examining names, sources, types, update frequency, geographic specificity, and alignment with the research on UK’s response to COVID-19.
4. Filters and Sorting Utilize filters to narrow results to recent, open-access datasets, applying sorting to prioritize by relevance or recency. Assess the precision of filtering and sorting mechanisms.
5. Dataset Actions Identify suitable datasets for longitudinal analysis of the pandemic’s effects, noting options for saving and sharing. Evaluate the efficiency and intuitiveness of these processes.
6. Feedback and Help Document any usability issues encountered and identify how the platform provides assistance. Explore the process for submitting feedback and seeking help within the platform.
3.4 Use of User Study Methods
3.4.1 Use of Concurrent Think-Aloud
This study uses moderate-complexity tasks, guided by [16]. Participants perform tasks that mix quick reactions with thoughtful analysis, such as choosing a relevant dataset and analysing its metadata. These tasks are more involved than basic queries but not as complex as detailed data analysis. This research employs a think-aloud technique as suggested by [7], giving real-time insight into participants’ thinking. VooV software captures both audio and visual data for comprehensive documentation.
3.4.2 Challenges in Concurrent Think-Aloud
One issue is that participants may feel awkward vocalising their thoughts, which can affect natural decision-making and skew results. Participants might also forget to verbalise their thoughts, especially when focused or performing difficult tasks. Another challenge is the risk of cognitive overload due to multitasking, especially in tasks requiring intense concentration [7].
3.4.3 Addressing Challenges During Concurrent Think-Aloud
@@
Used to help participants articulate their thoughts during this study. These are specific questions or statements aimed at prompting real-time reactions as participants use dataset search platforms. The goal is to get a detailed understanding of the user experience. Examples of thought prompts include: "What do you think of this dataset after selecting it?" and "Which filters are you using and why?".
3.4.4 Use of Stimulated Retrospective Think-Aloud
This research incorporates both Stimulated Retrospective Think-Aloud (SRTA) and Concurrent Think-Aloud (CTA) techniques to achieve a comprehensive view of user interactions [22, 7]. The SRTA method allows participants to reflect on their actions and thoughts, using visual aids or playback footage for reference. Structured SRTA questions guide this reflective process, creating a reliable setting for participants to share their experiences, thereby improving data quality. An example from this study illustrates this approach.
  • Retrospective Thought Prompt: "Reflecting on the search, do you remember your initial reactions? Did the results meet your expectations?"
  • Possible Retrospective Answer: "Initially, I was shocked by the sheer volume of information. I now realise my search terms were too broad, leading to the overload. If I were to do it again, I’d use more specific search criteria to narrow down the results."